stanbon 21772
having trouble to figure out how you solved the question about the baseball player hits a homerun on the average of once every 12 times at bat= would you mind letting me what to due
thanks

Answer by stanbon(75887)

Hope the additional explanation is helpful.
Let me know if there are statements you do
not understand.
---------------------------
A baseball player hits a homerun on the average of once every 12 times at bat. Use the idea of independent trials to answer the following questions.
---

If he bats 5 times in a game, what is the probability he will hit exactly one home run?
It's binomial with n=5, p = 1/12, x = 1
If you use a TI calculator you get binompdf(5,12^-1,1) = 0.2942
--------
If you do not have a TI: P(x=1) = 5C1(1/12)^1(11/12)^4
= 5*(1/12)(0.70607) = 0.2942
------------------------------------------
If he comes to bat 5 times, what is the probability he will hit at least one home run?
P(at least one) = 1 - P(no home runs) = 1-binompdf(5,12^-1,0) = 0.3528
OR
P(at least one) = 1-P(no home runs) = 1 - 5C0(1/12)^0(11/12)^5
= 1 - 1*1* = 1-0.6472 = 0.3528
------------------------------------------
If he comes to bat 5 times in a game, what is the expected value for the number of home runs he will hit?
E(x) = np = 5(1/12) = 5/12
==============================================
Cheers,
Stan H.